OpenStudy (anonymous):

You are camping at the Grand Canyon. When you pitch your tent at 1:00 P. M. the temperature is 81° F. When you go to bed at 11 P.M. the temperature is 61° F. What is the average rate of change in the temperature?

OpenStudy (anonymous):

\[\frac{61-81}{11-1}=\frac{-20}{10}=-2\]

OpenStudy (anonymous):

x - a represents the change in the input of the function f f(x) - f(a) represents the change in the function f as the input changes from a to x then the average rate of change A(x) is\[A(x)=\frac{f(x)-f(a)}{x-a}\]